# DiPT: Enhancing LLM Reasoning through Diversified Perspective-Taking

### v1

@@ Abstract
Keywords Reasoning  $\cdot$ Perspective-Taking  $\cdot$ LLM
@@ 2 Related Work
Numerous single-prompt (0-shot) methods have emerged to improve the model’s reasoning capabilities. One such method is (automatic/0-shot) chain-of-thought (CoT) Wei et al. (2022); Kojima et al. (2022), which instructs the model to provide a step-by-step explanation of the answer. This can be achieved by either incorporating examples with such explanations or by introducing an additional sentence in the prompt, "Let’s think Step by Step." Plan-and-solve (PS) method Wang et al. (2023a) is the extension of the CoT reasoning, which asks the model to first come up with the plan before solving the problem in a step-by-step manner. Recent work also derives theoretical analysis Feng et al. (2023) explaining how the transformers with chain-of-thought reasoning can solve mathematical problems that otherwise would not be possible without outputting the reasoning by the model. Another line of work Mekala et al. (2023); Deng et al. (2023) attempts to involve the model to simplify the query before actually solving the problem by asking the model to rephrase the query in the model’s simplified language. With a simplified query, the model can better understand the problem and proceed to solve the task. Analogical reasoners Yasunaga et al. (2024), on the other hand, instructs the LLM to self-generate similar examples to the query as demonstrations and then solve the problem. Overall, the common limitation of these methods is that they do not regulate how reasoning should be performed and, by default, adopt a single solution path. This can be attributed to various factors, such as the simplicity and computational efficiency of generating a single solution path, the lack of explicit rewards for diversity in the reasoning process in current evaluation metrics, and the assumption that a correct solution path indicates sufficient problem understanding. While investigating the mechanisms that encourage the generation of a single solution path is beyond the scope of this paper, we focus on studying the empirical benefits of incorporating multiple solution paths for both inference and training stages. Improving the effectiveness of single prompting naturally involves incorporating multiple prompts, such as CoT self-consistency Wang et al. (2022), least-to-most prompting Zhou et al. (2022), (probabilistic) tree-of-thoughts (ToT) Yao et al. (2023); Cao et al. (2023), or graph-of-thoughts Besta et al. (2023). These methods enhance responses by leveraging diverse model outputs. While diversified perspective-taking shows promise in improving reasoning based on multiple prompts by increasing the accuracy of individual prompts, this paper focuses on integrating diverse perspectives into zero-shot methods as a proof of concept.
@@ 3.1 DiPT as an Inference-Time Reasoning Enhancement Tool
For multiple-choice problems, DiPT prompts the model to re-think each option as a potential solution. In doing so, the model is encouraged to thoroughly consider each choice, rather than simply selecting the first most likely option based on next-word prediction, which might be subject to spurious correlations existing in the training data. For instance, consider the following text: “The backyard battles you staged with your green plastic army men were more exciting and almost certainly made more sense.” GPT-4-0613 labeled this text as having a positive sentiment, potentially due to the presence of certain words such as “exciting” and “make more sense,” which often appear in positive reviews. However, by explicitly prompting the model to think from the perspective that this review might be negative, it can better grasp the text’s negative tone and correct its answer. This example demonstrates how DiPT can help the model overcome misleading cues and better understand the overall context to make more accurate predictions. We refer the reader to Appendix A.5 and Appendix B.2 for detailed prompts and response.
For open-ended questions or free-text generation problems, DiPT prompts the model to consider different methods or approaches to solve the problem. For example, the question “What is the sum of all numbers between $-27\leq x<27$?” is surprisingly difficult for existing LLMs. Even when leveraging CoT, they often arrive at the wrong answer. On the other hand, with DiPT, the model would first generate potential methods to solve this question, such as the “arithmetic series method,” “symmetry method,” and “direct summation method” (see details in Figure 3). We provide the comparison of each method’s prompts in Figure 2. While some of these methods may still lead to incorrect answers, others would guide the model to the correct solution. By considering multiple approaches, the model can reflect upon its decisions and ultimately choose the correct answer. This example highlights the potential of DiPT to improve the model’s problem-solving capabilities, even for challenging questions that existing LLMs struggle with.
@@ 4 Experiment
This section presents experiments designed to investigate the following questions: (1) How does the integration of perspective-taking into existing reasoning methods impact their performance across various tasks? We evaluate its effect on both the accuracy and robustness to paraphrased problem statements. (Section 4.1) (2) What novel applications can be developed to harness DiPT’s advanced context understanding and accurate reasoning capabilities? Specifically, we will explore its potential in harmful query moderation and dataset error detection. (Section 4.2). (3) How does fine-tuning models on datasets enriched with perspective-taking affect their performance on both in-distribution and out-of-distribution tasks? (Section 4.3).
By addressing these questions, we aim to provide a comprehensive evaluation of the proposed approach and offer insights into its effectiveness, versatility, and generalizability.
@@ 4.1 DiPT Integration Impact on Inference
To understand the impact of perspective-taking on reasoning, we demonstrate the effect of adding DiPT to diverse reasoning methods. We considered three existing methods: CoT, which performs step-by-step reasoning; Rephrase and Respond (RaR), which rephrases and expands the question; and Analogical Reasoners (ANL), which self-generates examples similar to the problem. This diverse set allows us to assess the generalizability of DiPT across different reasoning paradigms. We emphasize that the goal is not to exhaustively evaluate DiPT with every state-of-the-art method. Instead, our focus is to understand the specific impact of perspective-taking on reasoning performance.
@@
We perform inference-stage experiments on $7$ tasks: AG News, CosmosQA, RTE, SST-5, SVAMP, TREC, and TruthfulQA. For AG News, SST-5, and TREC, we measure the Top-2 accuracy, as it is possible for an example to belong to multiple classes. For all other tasks, we apply Top-1 accuracy. We refer the reader to Appendix A for further details on datasets. We evaluate performance over $300$ test examples and report the average after $3$ runs. In the main paper, we report results on the GPT-4-Turbo (November) model Achiam et al. (2023), while we provide results on the open-weight model, Mistral7B-Instruct-v0.1 Jiang et al. (2023), in Appendix B.3. Additionally, we provide each DiPT prompt in Appendix A.4. We report 00-shot results of the target model with standard prompting; for each reasoning method, we report results when prompting with and without perspective-taking (DiPT+<Method-Name>) and the difference in the performance ($\Delta$).
@@
Standard (0-Shot) $79$ $83$ $87$ $86$
Chain-of-Thought $79$ $85$ $88$ $87$
DiPT+ Chain-of-Thought $82$ $89$ $91$ $93$
@@
Rephrase and Respond $80$ $83$ $89$ $89$
DiPT+ Rephrase and Respond $83$ $85$ $90$ $94$
@@
Analogical Reasoning $81$ $84$ $90$ $90$
DiPT+ Analogical Reasoning $84$ $88$ $90$ $94$
@@
In Table 1, we observe that adding perspective-taking to each of the reasoning methods improves performance in most cases with even $6\%$ increase for CoT in the TREC dataset. We observe performance increases for all cases except for the analogical reasoning with the RTE dataset, where the performance might have reached its peak due to potential labeling errors within the dataset. We will analyze these errors in detail in Section 4.2.
Figure 3: The comparison of responses between Standard, CoT, and DiPT+ CoT prompting.
To gain deeper insights into the positive quantitative results, Figure 3 presents an illustrative example. This example showcases how explicit exploration of multiple solution paths, enabled by DiPT in conjunction with CoT prompting, allows the language model to self-correct. Standard prompting and CoT prompting typically guide the model along a single path, increasing its susceptibility to errors (as seen in Figure 3, where the answers following their corresponding solution paths are incorrect). Conversely, DiPT prompts the model to explore alternative solutions. This capability allows for robust analysis and comparison of answers, ultimately leading the model to identify and correct errors, resulting in a correct final answer (shown in full in Appendix B.4).
@@
While current reasoning methods enhance the model’s capabilities, they may generate erroneous reasoning steps across various problem formulations, as noted in studies by Wang et al. (2023b); Lanham et al. (2023); Turpin et al. (2024). We examine whether incorporating perspective-taking into existing methods can enhance stability across different problem paraphrases, thus improving method reliability. To assess this, we evaluate each method’s output stability by measuring its sensitivity to paraphrased prompts. Specifically, we generate five paraphrases of the same queries used in Table 1 and report the mean performance across these iterations. Paraphrasing templates and examples are provided in Appendix A.6. Due to the automatic nature of paraphrasing, a few cases have lost their original meaning due to simplistic rephrasing, resulting in decreased performance across most scenarios in Table 2. However, we observe that all tested methods (CoT, RAR, and ANL) benefit from incorporating perspective-taking. This is evident in two key findings. First, across all methods, incorporating perspective-taking leads to the best overall performance on paraphrased problems. Second, the performance drops for methods with perspective-taking are usually smaller than those without it.
@@
4.2 Applications of DiPT Integration
Safety Moderation.
@@ 4.2 Applications of DiPT Integration
Dataset Labeling Error Detection.
We observe that although our method achieves high performance, it cannot reach $100\%$. Upon closer examination, we identified errors within datasets that prevented our method from achieving a perfect score. Current works in reasoning often use datasets commonly employed in NLP. However, for a dataset to serve as a reliable benchmark, it must exhibit high quality without errors. Otherwise, achieving high performance on inaccurately labeled data can mislead the comprehension of the method. Consequently, we aim to prevent such errors in these datasets. We apply our method to verify the labeling of these datasets and identify potential errors in the misalignment of the labels. Specifically, we employ DiPT on the gpt-4-1106-preview model to identify mismatched labels between the predicted and annotated labels. Then, to evaluate the correctness of error identification by our method, we leverage expertise evaluations from several powerful LLMs, including Bard/Gemini and Claude, in conjunction with our judgments. We use fine-grained metrics to better categorize the labeling errors: Wrong label, where all experts disagree with the original ground truth label, Ambiguous examples, where some experts disagree with the original label, and False Positives, where all experts agree with the original label.
SST-5
AG News
TREC
DBPedia
CosmosQA
SVAMP
TruthfulQA
RTE
Wrong 15 4 5 1 5 5 8 3
Ambiguous 4 0 2 0 6 0 3 5
False Positive 1 0 2 0 2 0 10 2
@@
XWinograd Test Multilingualism
Base Model 81.46 81.43 79.52 79.68 60.21 62.52 60.74 62.65
XWinograd
Plain 3K
81.43 81.36 81.95 79.46 60.26 62.82 61.53 62.56
XWinograd
CoT 3K
81.50 81.39 81.48 80.00 60.37 63.26 61.63 62.72
XWinograd
DiPT+ CoT 3K
81.50 81.41 82.15 80.06 60.47 63.35 61.90 62.78
Mistral7B
Base
Llama3-8B
Base
Mistral7B
Instruct-v0.2
Lamma3-8B
Instruct
Mistral7B
Base
Lamma3-8B
Mistral7B
Instruct-v0.2
Lamma3-8B
Instruct
Table 5: The fine-tuning results of four different models. The models are trained separately on OpenbookQA, GSM8K, CoQA, and XWinograd and evaluated on their test split (Left: in distribution) and on the associated domain (Right: in domain). Bold means the highest performance, and underlined means the second highest.

### v2

@@ Abstract
DiPT: Enhancing LLM Reasoning through Diversified Perspective-Taking
Hoang Anh Just Virginia Tech just@vt.edu                        Mahavir Dabas Virginia Tech mahavirdabas18@vt.edu                        Lifu Huang UC Davis lfuhuang@ucdavis.edu
Ming Jin Virginia Tech jinming@vt.edu                        Ruoxi Jia Virginia Tech ruoxijia@vt.edu
@@ 2 Related Work
Numerous single-prompt (0-shot) methods have emerged to improve the model’s reasoning capabilities. One such method is (automatic/0-shot) chain-of-thought (CoT) Wei et al. (2022); Kojima et al. (2022), which instructs the model to provide a step-by-step explanation of the answer. This can be achieved by either incorporating examples with such explanations or by introducing an additional sentence in the prompt, "Let’s think Step by Step." Plan-and-solve (PS) method Wang et al. (2023b) is the extension of the CoT reasoning, which asks the model to first come up with the plan before solving the problem in a step-by-step manner. Recent work also derives theoretical analysis Feng et al. (2023) explaining how the transformers with chain-of-thought reasoning can solve mathematical problems that otherwise would not be possible without outputting the reasoning by the model. Another line of work Mekala et al. (2023); Deng et al. (2023) attempts to involve the model to simplify the query before actually solving the problem by asking the model to rephrase the query in the model’s simplified language. With a simplified query, the model can better understand the problem and proceed to solve the task. Analogical reasoners Yasunaga et al. (2024), on the other hand, instructs the LLM to self-generate similar examples to the query as demonstrations and then solve the problem. Overall, the common limitation of these methods is that they do not regulate how reasoning should be performed and, by default, adopt a single solution path. This can be attributed to various factors, such as the simplicity and computational efficiency of generating a single solution path, the lack of explicit rewards for diversity in the reasoning process in current evaluation metrics, and the assumption that a correct solution path indicates sufficient problem understanding. While investigating the mechanisms that encourage the generation of a single solution path is beyond the scope of this paper, we focus on studying the empirical benefits of incorporating multiple solution paths for both inference and training stages.
Improving the effectiveness of single prompting naturally involves incorporating multiple prompts, such as CoT self-consistency Wang et al. (2022), least-to-most prompting Zhou et al. (2022), (probabilistic) tree-of-thoughts (ToT) Yao et al. (2023); Cao et al. (2023), or graph-of-thoughts Besta et al. (2023). These methods enhance responses by leveraging diverse model outputs. While diversified perspective-taking shows promise in improving reasoning based on multiple prompts by increasing the accuracy of individual prompts, this paper focuses on integrating diverse perspectives into zero-shot methods as a proof of concept. Concurrent with our work, perspective-taking has been effectively implemented to mitigate toxicity and bias in language models. By considering diverse audience perspectives, models can self-correct and reduce biases in their outputs (Xu et al., 2024). Wang et al. (2024) focuses on mitigating bias caused by false information in the prompt. However, it does not address improving reasoning or correcting false reasoning during generation.
@@ 3.1 DiPT as an Inference-Time Reasoning Enhancement Tool
For multiple-choice problems, DiPT prompts the model to re-think each option as a potential solution, i.e. prompting "Before choosing the answer, for each option explain if it is possible or not." In doing so, the model is encouraged to thoroughly consider each choice, rather than simply selecting the first most likely option based on next-word prediction, which might be subject to spurious correlations existing in the training data. For instance, consider the following text: “The backyard battles you staged with your green plastic army men were more exciting and almost certainly made more sense.” GPT-4-0613 labeled this text as having a positive sentiment, potentially due to the presence of certain words such as “exciting” and “make more sense,” which often appear in positive reviews. However, by explicitly prompting the model to think from the perspective that this review might be negative, it can better grasp the text’s negative tone and correct its answer. This example demonstrates how DiPT can help the model overcome misleading cues and better understand the overall context to make more accurate predictions. We refer the reader to Appendix A.5 and Appendix B.2 for detailed prompts and response.
For open-ended questions or free-text generation problems, DiPT prompts the model to consider different methods or approaches to solve the problem. For example, the question “What is the sum of all numbers between $-27\leq x<27$?” is surprisingly difficult for existing LLMs. Even when leveraging CoT, they often arrive at the wrong answer. On the other hand, with DiPT, the model would first generate potential methods to solve this question, such as the “arithmetic series method,” “symmetry method,” and “direct summation method”. Please see details in Figure 2 with the comparison of each method’s prompts. While some of these methods may still lead to incorrect answers, others would guide the model to the correct solution. By considering multiple approaches, the model can reflect upon its decisions and ultimately choose the correct answer. This example highlights the potential of DiPT to improve the model’s problem-solving capabilities, even for challenging questions that existing LLMs struggle with.
@@
While in the vanilla implementation, DiPT does not explicitly determine the number of perspectives to explore nor the strategy to choose the final decision, we explore them in Section 4.
@@ 4 Experiment
This section presents experiments designed to investigate the following questions: (1) How does the integration of perspective-taking into existing reasoning methods impact their performance across various tasks? We evaluate its effect on both the accuracy and robustness to paraphrased problem statements. (Section 4.1) (2) What novel applications can be developed to harness DiPT’s advanced context understanding and accurate reasoning capabilities? Specifically, we will explore its potential in harmful query moderation (Section 4.2) and dataset error detection (App B). (3) How does fine-tuning models on datasets enriched with perspective-taking affect their performance on both in-distribution and out-of-distribution tasks? (Section 4.3). By addressing these questions, we aim to provide a comprehensive evaluation of the proposed approach and offer insights into its effectiveness, versatility, and generalizability.
@@ 4.1 DiPT Integration Impact on Inference
To understand the impact of perspective-taking on reasoning, we demonstrate the effect of adding DiPT to diverse reasoning methods. We considered four existing methods: CoT, which performs step-by-step reasoning; Rephrase and Respond (RaR), which rephrases and expands the question; Analogical Reasoners (ANL), which self-generates examples similar to the problem; and CoT-SC, which samples multiple CoT generations (5 in our experiments) and chooses the final answer with majority vote. This diverse set allows us to assess the generalizability of DiPT across different reasoning paradigms. We emphasize that the goal is not to exhaustively evaluate DiPT with every state-of-the-art method. Instead, our focus is to understand the specific impact of perspective-taking on reasoning performance.
@@
We perform inference-stage experiments on $7$ tasks: AG News, CosmosQA, RTE, SST-5, SVAMP, TREC, TruthfulQA, DROP, MATH, and GPQA. For AG News, SST-5, and TREC, we measure the Top-2 accuracy, as it is possible for an example to belong to multiple classes. For all other tasks, we apply Top-1 accuracy or Exact Matching. We refer the reader to Appendix A for further details on datasets. We evaluate performance over $300$ test examples and report the average after $3$ runs with standard deviation. In the main paper, we report results on the GPT-4-Turbo (November) model Achiam et al. (2023), while we provide results on the open-weight model, Mistral7B-Instruct-v0.1 Jiang et al. (2023), in Appendix B.3. Additionally, we provide each DiPT prompt in Appendix A.4. We report 00-shot results of the target model with standard prompting; for each reasoning method, we report results when prompting with and without perspective-taking (DiPT+<Method-Name>) and the difference in the performance ($\Delta$).
@@
Standard (0-Shot) $79_{\pm 0.8}$ $83_{\pm 1.6}$ $87_{\pm 0.8}$ $86_{\pm 0.0}$
Chain-of-Thought $79_{\pm 0.8}$ $85_{\pm 1.6}$ $88_{\pm 0.9}$ $87_{\pm 0.0}$
DiPT+ Chain-of-Thought $82_{\pm 0.4}$ $89_{\pm 0.8}$ $91_{\pm 0.0}$ $93_{\pm 0.0}$
@@
Rephrase and Respond $80_{\pm 0.8}$ $83_{\pm 0.4}$ $89_{\pm 0.8}$ $89_{\pm 1.6}$
DiPT+ Rephrase and Respond $83_{\pm 0.8}$ $85_{\pm 0.8}$ $90_{\pm 0.5}$ $94_{\pm 0.0}$
@@
Analogical Reasoning $81_{\pm 2.4}$ $84_{\pm 1.2}$ $90_{\pm 0.0}$ $90_{\pm 0.0}$
DiPT+ Analogical Reasoning $84_{\pm 0.8}$ $88_{\pm 1.6}$ $90_{\pm 0.0}$ $94_{\pm 0.0}$
@@
Standard (0-Shot) CoT DiPT+ CoT $\Delta$ Performance
DROP $84_{\pm 1.6}$ $85_{\pm 0.9}$ $87_{\pm 0.4}$ $\uparrow 2$
MATH $86_{\pm 0.8}$ $88_{\pm 0.5}$ $90_{\pm 0.5}$ $\uparrow 3$
Table 2: Performance comparison between standard prompting, CoT and DiPT+CoT. Performance on free generation datasets.
CoT-SC DiPT+ CoT DiPT+ CoT-SC $\Delta$ Performance
DROP $85_{\pm 0.8}$ $87_{\pm 0.4}$ $88_{\pm 0.8}$ $\uparrow 3$
GPQA $60_{\pm 0.8}$ $60_{\pm 0.5}$ $62_{\pm 0.8}$ $\uparrow 2$
Table 3: Performance comparison with CoT-SC. Delta performance between CoT-SC and DiPT+CoT-SC.
@@
In Table 1, we observe that adding perspective-taking to each of the reasoning methods improves performance in most cases with even $6\%$ increase for CoT in the TREC dataset. We observe performance increases for all cases except for the analogical reasoning with the RTE dataset, where the performance might have reached its peak due to potential labeling errors within the dataset. We will analyze these errors in detail in Section B.
Additionally, we also observe in Table 2, that DiPT enhances CoT by improving performance by at least $2\%$ on both datasets. To gain deeper insights into the positive quantitative results, Figure 2 presents an illustrative example. This example showcases how explicit exploration of multiple solution paths, enabled by DiPT in conjunction with CoT prompting, allows the language model to self-correct. Standard prompting and CoT prompting typically guide the model along a single path, increasing its susceptibility to errors, where the answers following their corresponding solution paths are incorrect). Conversely, DiPT prompts the model to explore alternative solutions. This capability allows for robust analysis and comparison of answers, ultimately leading the model to identify and correct errors, resulting in a correct final answer (shown in full in Appendix B.4). While CoT-SC generates independent reasoning paths, they are not guaranteed to be coming from different perspectives. With this in mind, by enhancing CoT-SC explicitly with perspective-taking in each CoT generation, we observe in Table 3 that DiPT+CoT-SC can improve performance by even $3\%$. In sum, adding multiple perspectives to reasoning methods can help the model to arrive at accurate solutions more frequently. This approach mitigates potential wrong reasoning paths, ultimately improving the model’s overall performance.
@@
While current reasoning methods enhance the model’s capabilities, they may generate erroneous reasoning steps across various problem formulations, as noted in studies by Wang et al. (2023a); Lanham et al. (2023); Turpin et al. (2024). We examine whether incorporating perspective-taking into existing methods can enhance stability across different problem paraphrases, thus improving method reliability. To assess this, we evaluate each method’s output stability by measuring its sensitivity to paraphrased prompts. Specifically, we generate five paraphrases of the same queries used in Table 1 and report the mean performance across these iterations. Paraphrasing templates and examples are provided in Appendix A.6. Due to the automatic nature of paraphrasing, a few cases have lost their original meaning due to simplistic rephrasing, resulting in decreased performance across most scenarios in Table 4. However, we observe that all tested methods (CoT, RAR, and ANL) benefit from incorporating perspective-taking. This is evident in two key findings. First, across all methods, incorporating perspective-taking leads to the best overall performance on paraphrased problems. Second, the performance drops for methods with perspective-taking are usually smaller than those without it.
@@
Figure 3: The runtime plot of actual model output generation using DiPT+CoT prompting with a varying number of perspectives. The dotted line is the expected time projected from a single perspective. (Average time in seconds.)
Runtime Cost Analysis.
DiPT inherently encourages the model to generate multiple perspectives on a problem, potentially increasing overall generation. While the time to generate K perspectives might be expected to scale linearly (K times the time required for a single perspective), Figure 3 shows that the actual time scales sublinearly. This suggests that DiPT does not incur the expected costs as the number of perspectives increases. One possible explanation is that since all reasoning occurs simultaneously, one perspective can influence subsequent ones, causing the generation to converge on an answer more quickly, regardless of the correctness of the conclusion. Additionally, we note that for CoT-SC, the time does scale linearly with the number of paths.
Final Decision Strategies.
While the default implementation of DiPT  lets the model self-decide on the final decision based on results from multiple perspectives (and often resorts to majority voting), we now study different strategies for choosing the final answer: Repeat Decision, where we ask to repeat the answer, Verify Decision, where we ask to verify the decision given the analysis from different perspectives Dhuliawala et al. (2023), and Condition Consistency, where we ask to carefully check the satisfaction of problem conditions Weng et al. (2023) for choosing the final answer.
CosmosQA TruthfulQA RTE TREC
Default 82 89 91 93
[…]
Verify Decision 84 91 91 94
Condition Consistency 82 86 90 96
Table 5: Results on different decision strategies for DiPT+CoT.
In Table 5, we observe that Verify Decision consistently outperforms the default method for each considered dataset. Furthermore, several other methods demonstrate improvements over the default in several cases. These suggest that a principled, systematic way of choosing the final answer can improve upon the default, implicit way of choosing the final answer.
4.2 Applications: Safety Moderation.
@@
Table 7: The fine-tuning results of four different models. The models are trained separately on OpenbookQA, GSM8K, and CoQA and evaluated on their test split (Left: in distribution) and on the associated domain (Right: in domain). Bold means the highest performance, and underlined means the second highest.